Could someone help with 4B? I would like a fresh new answer please.4. (10 pts.) Assume two changes to the fast food industry: 1) a new machine that allows customers to make their own orders at
Answer #1

42 Production function is Q=25k-lok²_K2+62-0.22² 62 Now, marginal product of labor (MPL) 22- -*+6-0:42 and Marginal product oNOWO 2S -k 76-0.42 = 21 25-20k-2 - To Now substituting k=1 we qet! - » -146-0142 25=120217-2 - - => 5o me can 15- 12L = 25-52Now, total revenue = PxQ = $( 5 X 26.6) E $33 Profit will be Now, Total cosa = To tal revime - = $ (133-80) = $ 53 Scanned wi